A Complete Bibliography of Publications of Marsha J. Berger

Marsha J. Berger

Courant Institute - New York University 251 Mercer St. New York, NY

USA

Tel: +1 212 998 3305 FAX: +1 212 995 4121

E-mail: berger@cims.nyu.edu (Internet)

01 October 2024 Version 2.27

Abstract

This bibliography records publications of Marsha J. Berger.

Title word cross-reference

1 [BHOR83, HRBO83]. 3 [MBAW95, MAB03a]. 6 [MAB03b]. H [BHL03, AB02, BH02, BH12].

-Box [BHL03, BH02, BH12]. -D [MAB03a]. -Dimensional [BHOR83, HRBO83]. -DOF [MAB03b]. -Refinement [AB02].

10th [Ame83]. 12th [DG95]. 18 [MAB04b]. 1983 [BCF83]. 1984 [SB85]. 1985 [DeG85]. 1989 [Ano89]. 1990 [EG90]. 1993 [MMM93]. 1996 [KFC97, SK96].

2 [BS94]. 2005 [ACM05].

3D [BME93].

50th [Gau94].

70th [BW05].

9th [Ano89].

A-18 [MAB04b]. Accuracy [BAM96, BM95, BM96]. accurate [KB06]. ACM [ACM05]. ACM/IEEE [ACM05]. across [BB85]. Adaptation [AMB95]. Adaptive [ABM98a, AB02, BBSW94, Ber82, BHOR83, Ber83, BJ84, BO84, BJ85a, Ber85a, Ber86, Ber87a, BL89, Ber89a, BC89, Ber89b, BL90a, BS93, BAM96, BL98, BL24, HRBO83, WB96, WKB^{+}08, ZPCL08, BJ85b, BL91a, BCHL09, BL23, RPB99, BCF83]. adaptively [ABA95, ABA00, LGB11]. Administration [MMM93]. Advanced [ZPCL08]. Aeronautics [ABB^{+}06, MMM93]. Aerospace [MAB05, MAB07]. AFOSR [SK96]. Agglomeration [BVM93]. AIAA [DG95]. Airborne [WKB^{+}08]. Airframes [MAB02, MAB04a]. al [MMM93]. Algorithm [BL89, BR92, BG24, BK88, BG21, GAB^{+}22]. Algorithms [BL98]. Aligned [MB13, BH02]. AMR [BS94]. Analysis [BB^{+}78, BAM05]. Anniversary [Gau94]. Antony [BW05]. Applications [ABM04, ABA06, ABB^{+}06, Ber89b, EG90, MAB05, MAB07, MBAW95, ZPCL08, HK02, S^{+}83]. Approach [ABA06]. Approaches [MAB03a]. Approximation [BHL03]. April [MMM93]. Arbitrary [BL89]. aspect [BA98]. Aspects [BA98]. Asteroid [Ber22, BL23]. Asteroid-generated [Ber22]. Aug [Ame83]. August [DeG85, Gau94]. Automatic [BJ85a, BME93, DAB^{+}99, WB96].

based [ABM97, ABM98b]. Benchmarks [ABB^{+}06]. birthday [BW05]. blowing [BK88]. blowing-up [BK88]. Boundaries [AB02, MAB03a, ABA95, ABA00]. Boundary [ABA06, BL90b, MB17, BH12, GAB^{+}22, HB14, KB06, MB14, RPB99]. Boundary-Layer [ABA06]. Box [BHL03, BH02, BH12]. British [Gau94]. Brunswick [SK96]. Buffalo [Ano89]. building [MAB^{+}16].

CA [KFC97, DG95]. calculation [BK88]. Calculations [BL90b]. Canada [Ame83]. Cartesian [DAB^{+}99, AMB95, ABA95, ABM97, ABM98a, ABM98b, ABA00, AB02, ABM04, ABA06, BL89, BL90a, BL90b, BL91a, BL91b, BME93, BM95, BM96, BA98, BAA01, Ber02, FB99, HB14, MB14, MBAW95, MAB02, MAB03a, MAB03b, MAB04a, MAB04b, WB96, WKB^{+}08]. Cartesian/Quad [WB96]. celebration [BW05]. cell [Ber15, BG21]. Cells [MB17, Ber15, Ber17]. century [Gau94]. CFD [SK96, ABM04, BAMM05]. Characterization [ABB^{+}06]. Circulation [ASB75]. Clawpack [MAB^{+}16]. Clustering [BR92]. CM [BS94]. CM-2 [BS94]. collection [DG95]. College [BCF83]. Colorado [MMM93]. Columbia [MAB05, MAB07, ABB^{+}06, Gau94]. Complex [ABA06, BL91b, BAM96, DAB^{+}99, BM95, BM96, BAA01, BH02, Ber04, FB99, HK02]. component [ABM97, ABM98b]. component-based [ABM97, ABM98b]. computation [Ame83, Gau94]. Computational [Ano89, DG95, ZPCL08, BCF83, Gau94]. computations [EOS85]. Computing [S^{+}93, S^{+}83]. Conditions [BL90b, KB06]. Conference [Ano89, DeG85, EG90, KFC97, MMM93, SK96, S^{+}93, SB85, ACM05, DG95]. Configurations [ABA06]. congress [Ame83]. Connection [BS93]. Conservation [BBSW94, Ber87b, BHL03, HBL05, BH02]. Coordinate [MB13]. Copper [MMM93]. Corrector [BOR81]. cosponsored [MMM93]. curved [KB06]. Curves [ABM04]. Cut [Ber17, MB17, Ber15, BG21].

D [MBAW95, MAB03a]. Data

[Ber83, Ber86]. Dependent [BOR81]. Detailed [ABB^{+}06]. Diego [DG95]. Difference [Ber87a, BL91b]. Differential [Ber82, BO84, BL90a, BCF83, BL91a]. Dimensional [BBSW94, BHOR83, MB13, BLS94, HRBO83]. Dispersion [WKB^{+}08]. Dispersive [BL24, BL23]. Distributed [ZPCL08]. DOF [MAB03b]. Dynamic [SK96]. Dynamics [Ano89, BHOR83, Ber87a, KFC97, SB85, DG95, HRBO83].

ecosystem [MAB^{+}16]. efficient [ABM97, ABM98b]. Elliptic [Sch81].

Embedded [AB02, HBL05, MB17, ABA95, ABA00, BH12, GAB^{+}22, HB14, MB14]. engineering [Ame83]. Equations [Ber82, BJ84, BO84, BJ85a, Ber85a, BL89, BL90a, BCF83, BJ85b, BL91a]. Error [AB02]. Estimation [AB02]. Euler [BJ84, BJ85a, BJ85b, BL89, BME93, MBAW95]. Explicit [MB17, MB14].

F [MAB04b]. F/A [MAB04b]. F/A-18 [MAB04b]. Fifteenth [KFC97]. Filling [ABM04]. Finite [Ber87a, BLS94, HK02, MB13, BCHL09, BG21, LGB11, HK02]. First [SK96]. Flow [BAMM05, FB99, WB96, BM95, BM96, Ber04]. Flows [BME93, DAB^{+}99]. Fluid [Ano89, Ber87a, KFC97, SB85, DG95, EOS85]. France [HK02, SB85].

Gas [BHOR83, HRBO83]. generated [Ber22]. Generation [ABM98a, Ber86, BR92, BME93, DAB^{+}99, ABM97, ABM98b, Ber02, Ber04]. Geometries [BL89, BL91b, DAB^{+}99, HBL05, BM95, BM96, BH02, Ber04, FB99, KB06]. Geometry [BAM96, ABM97, ABM98b, BAA01, Ber02]. Gif [SB85]. Gif-sur-Yvette [SB85]. Gif-sur-Yvette/France [SB85]. Grid [BJ85a, Ber86, Ber87b, BL90b, BR92, BME93, BH02, DAB^{+}99, HBL05, MBAW95, WB96, BM95, BM96, Ber02, Ber04, HB14]. Grids [BL91b, BHL03, BAM05, ABA95, ABA00, BLS94, BH12, FB99, GAB^{+}22]. Guide [BB^{+}78].

half [Gau94]. half-century [Gau94]. held [KFC97, MMM93, S^{+}93]. High [DAB^{+}99, HBL05, KB06, MAB05, MAB07]. High-order [KB06]. High-Resolution [HBL05]. High-Reynolds [DAB^{+}99]. Hybrid [BAMM05, DAB^{+}99]. Hybrid-Cartesian [DAB^{+}99].

Hydrodynamics [BC89]. Hyperbolic [BBSW94, Ber82, BO84, Ber85a, BL90a, BL98, BHL03, EG90, BL91a, HB14, MAB^{+}16].

IEEE [ACM05]. III [HK02]. immersed [RPB99]. impact [BL23]. implementation [KB06]. Implicit [BL24, MB17, MAB03a, MB14]. Independent [WB96]. Infrastructures [ZPCL08]. Interfaces [Ber85b, Ber87b]. International [DeG85, EG90, HK02, KFC97, SB85]. involving [FB99]. Irregular [BHL03, BAM05, BLS94]. Island [HK02]. issue [BW05].

Jameson [BW05]. Jersey [SK96]. June [Ano89, DG95, EG90, KFC97, SK96].

Lake [ASB75]. Large [EOS85]. Large-scale [EOS85]. Laws [BBSW94, BHL03, HBL05, BH02]. Layer [ABA06]. Level [MAB02, MAB04a]. Library [BB^{+}78]. Limiters [BAM05, MB13]. Local [BC89]. Logically [BCHL09].

Machine [BS93]. March [S^{+}93].

mathematics [Gau94, S^{+}83, Gau94]. MD [BCF83]. mechanics [EOS85]. merging [Ber15]. Mesh [AMB95, ABM98a, ABA06, BBSW94, Ber82, BHOR83, Ber83, BO84, Ber85a, Ber85b, BL89, Ber89a, BC89, Ber89b, BL90a, BS93, BL98, BL24, ABM97, ABM98b, BA98, HRBO83]. Meshes [AB02, BL90a, MB13, MB17, BL91a, BAA01, Ber17, BG21, MB14]. Method [BJ84, HBL05, MBAW95, MAB02, MAB03a, MAB03b, MAB04a, MAB04b, ABA95, ABA00, BJ85b, BM95, BM96, BH12, RPB99]. Methods [AMB95, ABM04, BOR81, Ber87a, BAM96, BHL03, EG90, KFC97, MMM93, SB85, WKB^{+}08, BCF83, BLS94, BA98, BH02, BCHL09, HB14, LGB11]. Michigan [ASB75]. mixed [MB14]. Modeling [AMB95, WKB^{+}08, Ber02, Ame83]. modelling [LGB11]. Models [ASB75]. Monterey [KFC97]. Montreal [Ame83]. Motion [MAB03b, SK96]. Mountain [MMM93]. Moving [MAB03a, FB99]. Multi [MAB02, MAB04a]. Multi-Level [MAB02, MAB04a]. Multigrid [BJ84, BVM93, MMM93, BJ85b, BAA01]. Multilevel [AB02, ABA95, ABA00]. Multiprocessors [BB85, BB87].

NASA [MAB05, MAB07]. National [MMM93]. Ninth [SB85]. Non [BB87, MB13]. Non-Coordinate-Aligned [MB13]. Non-Uniform [BB87]. Norfolk [S^{+}93]. note [Ber15]. November [ACM05]. Number [DAB^{+}99]. Numerical [ASB75, BB^{+}78, EG90, KFC97, MAB02, MAB04a, SB85, BK88].

one [BLS94]. one-dimensional [BLS94]. open [MAB^{+}16]. order [KB06].

papers [DG95]. Parabolic [BOR81]. Paradigm [BAMM05]. Parallel [BOR81, Ber89a, BAA01, DeG85, S^{+}93, ZPCL08, ABA95, ABA00]. Park [BCF83]. Partial [Ber82, BO84, BL90a, BCF83, BL91a]. Partitioning [BB85, BB87]. PDEs [BB85, MAB^{+}16]. Performance [ABB^{+}06, BAMM05]. physical [S^{+}83]. pieces [Ber04]. Point [BR92]. Porquerolles [HK02]. Predictor [BOR81]. Predictor-Corrector [BOR81]. Preface [BW05]. problem [Sch81]. Problems [BOR81, BB87, EG90, HB14]. Proceedings [BCF83, EG90, KFC97, MMM93, SB85, ACM05, DeG85, SK96, S^{+}93]. Processing [DeG85, S^{+}93, Ber02]. Processors [BOR81, Ber89a]. Program [BB^{+}78]. Programming [BAMM05]. Propagation [BL98, BL24, BL23]. Provably [BG24]. Putting [Ber04].

Quad [WB96].

ratios [BA98]. rectangular [BCHL09]. Redistribution [BG24, BG21, GAB^{+}22]. refined [ABA95, ABA00, LGB11]. Refinement [AB02, BBSW94, Ber82, BHOR83, Ber83, BO84, BJ85a, Ber85a, Ber85b, Ber89a, BC89, Ber89b, BL90a, BS93, BL98, BL24, BL91a, BCHL09, HRBO83]. rescaling [BK88]. Resolution [HBL05, MAB05, MAB07]. review [Ber22]. Reynolds [DAB^{+}99]. Robust [ABM97, ABM98b]. Rolling [MAB02, MAB04a]. Rolling-Airframes [MAB02, MAB04a]. Rotated [BL91b, HBL05].

San [DG95]. scale [EOS85]. Scheme [BL91b, MB17, MB14]. Schemes [MB13, BG21]. sciences [S^{+}83]. Scientific [S^{+}93, Ame83, S^{+}83]. Seattle [ACM05]. Separation [MAB04b]. Shock [BC89]. SIAM [S^{+}93]. simplified [BH12]. Simulation [MAB02, MAB04a, Ame83]. Simulations [MAB03b, MAB04b, BL23, FB99]. Sixth [S^{+}93, MMM93]. Slope [BAM05, MB13]. solid [KB06]. Solution [BOR81]. Solutions [WB96, BK88]. Solver [BAMM05, WB96]. solvers [Ber04, Ber17, Sch81]. solving [MAB^{+}16]. Some [Ber89b]. source [MAB^{+}16]. Space [ABM04, MMM93]. Space-Filling [ABM04]. Special [BW05]. sphere [BCHL09]. Stability [Ber85b, Ber15]. Stable [BL90b, BG24]. State [BG24, BG21, GAB^{+}22]. steady [BM95, BM96]. stepping [MB14]. Store [MAB04b]. Strategy [BB85, BB87]. Structured [BS93]. Structures [Ber83, Ber86]. Supercomputer [MAB05, MAB07]. Supercomputing [ACM05]. Surface [AMB95, Ber02]. Sweden [EG90]. Symposium [Gau94, HK02]. Systems [Ame83, BL98].

technical [DG95]. test [BM95, BM96]. Theory [EG90]. Third [HK02, EG90]. Three [BBSW94]. Three-Dimensional [BBSW94]. Time [BOR81, MB14]. Time-Dependent [BOR81]. together [Ber04]. Triangle [Ber02]. Tsunami [BL24, LGB11, BL23]. tsunamis [Ber22]. Two [MB13]. Two-Dimensional [MB13].

Uniform [BB87]. Unstructured [BVM93]. Uppsala [EG90]. USA [BCF83, KFC97, S^{+}93]. User [BB^{+}78]. Using [BL98, MAB07, MAB02, MAB04a, ABB^{+}06, BAMM05, MAB05].

VA [S^{+}93]. Vancouver [Gau94]. Verification [ASB75]. version [RPB99]. Viscous [WB96]. Volume [MB13, BLS94, BCHL09, BG21, LGB11]. volumes [HK02].

WA [ACM05]. wall [KB06]. Wave [BL98]. Wave-Propagation [BL98]. Weighted [BG24, GAB^{+}22]. Workshop [BCF83, MMM93]. World [Ame83].

York [Ano89]. Yvette/France [SB85].

References

Aftosmis:2002:MEE

[AB02] M. J. Aftosmis and M. J.

Berger. Multilevel error estimation and adaptive h-refinement for Cartesian meshes with embedded boundaries. Technical Report AIAA 2002-0863, American Institute of Aeronautics and Astronautics, 1801 Alexander Bell Drive, Suite 500, Reston, VA 20191, USA, January 14–17, 2002. 14 pp.

URL http://people.nas.nasa.
gov/~aftosmis/publications/
aiaa2002-0863.pdf.

Aftosmis:1995:PMM

[ABA95] Michael Aftosmis, Marsha Berger, and Gedas Adomavicius. A parallel multilevel method for adaptively refined Cartesian grids with embedded boundaries. In Proceedings of the 12th AIAA Computational Fluid Dynamics Conference, San Diego, CA, June 1995, page ???? American Institute of Aeronautics and Astronautics, 1801 Alexander Bell Drive, Suite 500, Reston, VA 20191, USA, January 1995. AIAA 95-1725CP.

Aftosmis:2000:PMM

[ABA00] Michael Aftosmis, Marsha Berger, and Gedas Adomavicius. A parallel multilevel method for adaptively refined Cartesian grids with embedded boundaries. Technical Report AIAA 2000-0808, American Institute of Aeronautics and Astronautics, 1801 Alexander Bell Drive, Suite 500, Reston, VA 20191, USA, January 2000.

Aftosmis:2006:ACM

[ABA06] M. J. Aftosmis, M. J. Berger, and J. J. Alonso. Applications of a Cartesian mesh boundary-layer approach for complex configurations. AIAA Journal, page ??, January 2006.

Aftosmis:2006:DPC

[ABB^{+}06] M. Aftosmis, M. Berger, R. Biswas, M. J. Djomehri, R. Hood, H. Jin, and C. Kiris. A detailed performance characterization of Columbia using aeronautics benchmarks and applications. AIAA Journal, ??:??, January 2006. AIAA Paper 2006-84, Reno, NV.

Aftosmis:1997:REC

[ABM97] M. J. Aftosmis, M. J. Berger, and J. E. Melton. Robust and efficient Cartesian mesh generation for component-based geometry. Technical Report AIAA 97-0196, American Institute of Aeronautics and Astronautics, 1801 Alexander Bell Drive, Suite 500, Reston, VA 20191, USA, January 1997. Published in [ABM98b].

Aftosmis:1998:ACM

[ABM98a] M. J. Aftosmis, M. J. Berger, and J. E. Melton. Adaptive Cartesian mesh generation. In J. Thompson, B. Soni, and N. Weatherill, editors, Handbook of Grid Generation, page ?? CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1998. ISBN ???? LCCN ????

Aftosmis:1998:REC

[ABM98b] M. J. Aftosmis, M. J. Berger, and J. E. Melton. Robust and efficient Cartesian mesh generation for component-based geometry. AIAA Journal, 36(6):952– 960, June 1998. See report [ABM97].

Aftosmis:2004:ASF

[ABM04] M. J. Aftosmis, M. J. Berger, and S. M. Murman. Applications of space-filling curves to Cartesian methods for CFD. Technical Report AIAA 2004-1232, American Institute of Aeronautics and Astronautics, 1801 Alexander Bell Drive, Suite 500, Reston, VA 20191, USA, January 5–8, 2004. 12 pp.

                                        URL http://people.nas.nasa.
                                        gov/~aftosmis/publications/
                                        aiaa2004_1232.pdf.

ACM:2005:PAI

[ACM05] ACM, editor. Proceedings of the 2005 ACM/IEEE conference on Supercomputing 2005, Seattle, WA, November 12–18 2005. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2005. ISBN 1-59593-061-2. LCCN ????

Aftosmis:1995:ASM

[AMB95] M. Aftosmis, J. Melton, and M. Berger. Adaptation and surface modeling for Cartesian mesh methods. In Dolling and Grossman [DG95], page ?? ISBN ???? LCCN ???? AIAA Paper 95-1725.

Ames:1983:MSE

[Ame83] William F. Ames, editor. Modeling and simulation in engineering: Systems simulation and scientific computation (World congress) (10th: 1982 Aug: Montreal, Canada). North-Holland, Amsterdam, The Netherlands, 1983. ISBN 0-444-86606-X. LCCN ????

Anonymous:1989:CFD

[Ano89] Anonymous, editor. 9th Computational Fluid Dynamics Conference, Buffalo, New York, June, 1989. American Institute of Aeronautics and Astronautics, Aerospace Center, 370 L’Enfant Promenade, SW, Washington, DC 20024-2518, 1989. ISBN ???? LCCN ????

Allender:1975:VNC

[ASB75] J. Allender, K. Saunders, and M. Berger. Verification of numerical circulation models for Lake Michigan. Technical Report 43, Canada Centre for Inland Waters, Burlington, ON, Canada, 1975.

Berger:1998:AAR

[BA98] Marsha Berger and Michael Aftosmis. Aspects (and aspect ratios) of Cartesian mesh methods. In Sixteenth International Conference on Numerical Methods in Fluid Dynamics (Arcachon, 1998), volume 515 of Lecture Notes in Physics, pages 1– 12. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 3-540-65153-5.

Berger:2001:PMC

[BAA01] Marsha Berger, Michael Aftosmis, and Gedas Adomavicius. Parallel multigrid on Cartesian meshes with complex geometry. In Proceedings of the 8th International Congress on Parallel Computational Fluid Dynamics (Trondheim, Norway, June, 2000), pages 283–290. North-Holland, Amsterdam, The Netherlands, 2001. ISBN 0-444-50673-X.

Berger:1996:AAM

[BAM96] Marsha Berger, Michael Aftosmis, and John Melton. Accuracy, adaptive methods and complex geometry. In Sakell and Knight [SK96], page ?? ISBN ???? LCCN ????

Berger:2005:ASL

[BAM05] Marsha Berger, Michael J. Aftosmis, and Scott M. Murman. Analysis of slope limiters on irregular grids. AIAA Paper 2005-0490, American Institute for Aeronautics and Astronautics, Reno, NV, USA, May 2005. 22 pp.

                                        URL http://citeseerx.ist.
                                        psu.edu/viewdoc/summary?doi=
                                        10.1.1.139.8456;   http://
                                        www.nas.nasa.gov/News/Techreports/
                                        2005/PDF/nas-05-007.pdf. Also

NASA TM NAS-05-007.

Berger:2005:PNC

[BAMM05] M. J. Berger, M. J. Aftosmis, D. D. Marshall, and S. M. Murman. Performance of a new CFD flow solver using a hybrid programming paradigm. Journal of Parallel and Distributed Computing, 65(4):414–423, April 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Berger:1978:NAP

[BB^{+}78] M. Berger, J. Bolstad, et al. Numerical Analysis Program Library User’s Guide. Stanford Linear Accelerator Center Computing Services, Stanford, CA, USA, 1978.

Berger:1985:PSP

[BB85] Marsha Berger and Shahid Bokhari. A partitioning strategy for PDEs across multiprocessors. In DeGroot [DeG85], pages 166–170. ISBN 0-8186-0637-1 (paperback), 0-8186-8637-5 (hardback), 0-8186-4637-3 (microfiche). LCCN QA76.6 .I613 1985; QA76.6 .I548 1985.

Berger:1987:PSN

[BB87] Marsha Berger and Shahid Bokhari. A partitioning strategy for non-uniform problems on multiprocessors. IEEE Transactions on Computers, C-36(5):570– 580, May 1987. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). ICASE Report No. 85-55, Nov. 1985.

Bell:1994:TDA

[BBSW94] John Bell, Marsha Berger, Jeff Saltzman, and Mike Welcome. Three-dimensional adaptive mesh refinement for hyperbolic conservation laws. SIAM Journal on Scientific Computing, 15(1):127– 138, January 1994. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Berger:1989:LAM

[BC89] Marsha Berger and Phillip Colella. Local adaptive mesh refinement for shock hydrodynamics. Journal of computational physics, 82(1):64–84, May 1989.

CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). Lawrence Livermore Laboratory Report No. UCRL-97196.

Babuska:1983:ACM

[BCF83] Ivo Babuska, Jagdish Chandra, and Joseph E. Flaherty, editors. Adaptive computational methods for partial differential equations: Proceedings of the Workshop, College Park, MD, USA, 1983. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1983. ISBN 0-89871-191-6. LCCN QA377 .A29 1983.

Berger:2009:LRF

[BCHL09] Marsha J. Berger, Donna A. Calhoun, Christiane Helzel, and Randall J. LeVeque. Logically rectangular finite volume methods with adaptive refinement on the sphere. Philosophical Transactions of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 367(1907): 4483–4496, 2009. CODEN PTR-MAD, PTMSFB. ISSN 1364-503X (print), 1471-2962 (electronic).

Berger:1982:AMR

[Ber82] Marsha J. Berger. Adaptive Mesh Refinement for Hyperbolic Partial Differential Equations. Ph.D. dissertation, Department of Computer Science, Stanford University, Stanford, CA, USA, August 1982. vi + 118 pp. Computer Science Report No. STAN-CS-82-

Berger:1983:DSA

[Ber83] Marsha Berger. Data structures for adaptive mesh refinement. In Babuska et al. [BCF83], pages 237–251. ISBN 0-89871-191-6. LCCN QA377 .A29 1983.

Berger:1985:AMR

[Ber85a] Marsha Berger. Adaptive mesh refinement for hyperbolic equations. In Engquist et al. [EOS85], pages 31–40. ISBN 0-8218-1122-3. LCCN QA901 .L37 1985. Two volumes.

Berger:1985:SIM

[Ber85b] Marsha J. Berger. Stability of interfaces with mesh refinement. Mathematics of Computation, 45(172):301–318, October 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). ICASE Report No. 83-42, August, 1983.

Berger:1986:DSA

[Ber86] Marsha J. Berger. Data structures for adaptive grid generation. SIAM Journal on Scientific and Statistical Computing, 7(3):904– 916, July 1986. CODEN SIJCD4. ISSN 0196-5204.

Berger:1987:AFD

[Ber87a] Marsha Berger. Adaptive finite difference methods in fluid dynamics. NYU Report DOE/ER/03077-277, New York University, New York, NY, USA, February 1987. ???? pp. Lecture notes prepared for the short course in Computational Fluid Dynamics at the von Karman Institute for Fluid Dynamics, March 1987.

Berger:1987:CGI

[Ber87b] Marsha J. Berger. On conservation at grid interfaces. SIAM Journal on Numerical Analysis, 24(5):967–984, October 1987. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). ICASE Report No. 84-43, September, 1984.

Berger:1989:AMR

[Ber89a] Marsha Berger. Adaptive mesh refinement for parallel processors. In G. Rodrigue, editor, Parallel Processing for Scientific Computing, Proc. 1987 SIAM Conference on Parallel Processing, Los Angeles, California, page ?? Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1989. ISBN ?? LCCN ?? Ultracomputer Note 130, Dec. 1987.

Berger:1989:SAA

[Ber89b] Marsha Berger. Some applications of adaptive mesh refinement. In J. Flaherty et al., editors, Adaptive Methods for Partial Differential Equations: Proc. Workshop on Adaptive Computational Methods for Partial Differential Equations, RPI, Oct. 1988, pages 150–159. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1989. ISBN ?? LCCN ??

Berger:2002:VNC

[Ber02] M. Berger. Triangle geometry processing for surface modeling and Cartesian grid generation. U.S. Patent # 6445390, September 3, 2002.

Berger:2004:PTP

[Ber04] Marsha Berger. Putting together the pieces: grid generation and flow solvers for complex geometries. In Applied mathematics entering the 21st Century, volume 116 of SIAM Proceedings in Applied Mathematics, pages 25–35. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2004. ISBN 0-89871-559-8.

Berger:2015:NSC

[Ber15] Marsha Berger. A note on the stability of cut cells and cell merging. Applied Numerical Mathematics: Transactions of IMACS, 96(??): 180–186, October 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0168927415000896.

Berger:2017:CCM

[Ber17] M. Berger. Cut cells: meshes and solvers. In Handbook of numerical methods for hyperbolic problems, volume 18 of Handb. Numer. Anal., pages 1–22. Elsevier/ North-Holland, Amsterdam, The Netherlands, 2017. ISBN 0-444-63910-1.

Berger:2022:AGT

[Ber22]  Marsha Berger.  Asteroid-

generated tsunamis: a review. In Recent advances in industrial and applied mathematics, volume 1 of ICIAM 2019 SEMA SIMAI Springer Ser., pages 3– 17. Springer, Cham, Switzerland, 2022. ISBN 3-030-86235-6; 3-030-86236-4.

Berger:2021:SRA

[BG21] Marsha Berger and Andrew Giuliani. A state redistribution algorithm for finite volume schemes on cut cell meshes. Journal of computational physics, 428(??): Article 109820, March 1, 2021. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0021999120305945.

Berger:2024:NPS

[BG24] Marsha Berger and Andrew Giuliani. A new provably stable weighted state redistribution algorithm. SIAM Journal on Scientific Computing, 46(5):A2848– A2873, ???? 2024. CO-DEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

                                        URL https://epubs.siam.org/
                                        doi/10.1137/23M1597484.

Berger:2002:GAB

[BH02] Marsha J. Berger and Christiane Helzel. Grid aligned h-box methods for conservation laws in complex geometries. In Herbin and Kröner [HK02], pages 277– 284. ISBN 1-903996-34-1. LCCN QA911 .I56 2002.

Berger:2012:SBM

[BH12] Marsha Berger and Christiane Helzel. A simplified h-box method for embedded boundary grids. SIAM Journal on Scientific Computing, 34(2):A861– A888, ???? 2012. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Berger:2003:BMA

[BHL03] Marsha J. Berger, Christiane Helzel, and Randall J. LeVeque. H-box methods for the approximation of hyperbolic conservation laws on irregular grids. SIAM Journal on Numerical Analysis, 41(3):893–918, June 2003. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/40539.

Berger:1983:AMR

[BHOR83] Marsha Berger, G. Hedstrom, J. Oliger, and G. Rodrigue. Adaptive mesh refinement for 1dimensional gas dynamics. In Ames [Ame83], pages 43–47. ISBN 0-444-86606-X. LCCN ???? Also as Lawrence Livermore UCRL-87522.

Berger:1984:AMM

[BJ84] Marsha Berger and A. Jameson.

An adaptive multigrid method for the Euler equations. In Soubbaramayer and Boujot [SB85], page ?? ISBN 0-387-13917-6. LCCN QA911 .I54 1984; QA911 .I58 1984; TA357 .I564 1984. DM82.00 (West Germany).

Berger:1985:AAG

[BJ85a] Marsha Berger and A. Jameson. Automatic adaptive grid refinement for the Euler equations. AIAA Journal, 23(4):561– 568, April 1985. MAE Report No. 1633, October 1983, Princeton University, and NYU Report No. DOE/ER/03077-202, October 1983.

Berger:1985:AMM

[BJ85b] Marsha J. Berger and Antony Jameson. An adaptive multigrid method for the Euler equations. In Soubbaramayer and Boujot [SB85], pages 92–97. ISBN 0-387-13917-6. LCCN QA911 .I54 1984; QA911 .I58 1984; TA357 .I564 1984. DM82.00 (West Germany).

Berger:1988:RAN

[BK88] Marsha Berger and Robert V. Kohn. A rescaling algorithm for the numerical calculation of blowing-up solutions. Communications on Pure and Applied Mathematics (New York), 41(6): 841–863, 1988. CODEN CPA-MAT. ISSN 0010-3640 (print), 1097-0312 (electronic).

Berger:1989:ACM

[BL89] Marsha Berger and Randy Le-Veque. An adaptive Cartesian mesh algorithm for the Euler equations in arbitrary geometries. In Anonymous [Ano89], page ?? ISBN ???? LCCN ????

Berger:1990:CMA

[BL90a] Marsha Berger and Randy Le-Veque. Cartesian meshes and adaptive mesh refinement for hyperbolic partial differential equations. In Engquist and Gustafsson [EG90], page ?? ISBN 91-44-33421-4 (Studentlitteratur), 0-86238-285-8 (Chartwell Bratt). LCCN QA377.I563 1990. Two volumes.

Berger:1990:SBC

[BL90b] Marsha Berger and Randy Le-Veque. Stable boundary conditions for Cartesian grid calculations. In Proc. Symposium on Computational Technology for Flight Vehicles, page ?? Pergamon Press, ??, November 1990. ISBN ?? LCCN ?? ICASE Report No. 90-37, May, 1990.

Berger:1991:CMA

[BL91a] Marsha Berger and Randall Le-Veque. Cartesian meshes and adaptive refinement for hyperbolic partial differential equations. In Third International Conference on Hyperbolic Problems, Vol. I, II (Uppsala, 1990), pages 67–73. Studentlitteratur, Lund, Sweden, 1991.

Berger:1991:RDS

[BL91b] Marsha Berger and Randy Le-Veque. A rotated difference scheme for Cartesian grids in complex geometries. In AIAA 91-1602, 10th AIAA Computational Fluid Dynamics Conference, Honolulu, June, 1991, page ?? American Institute of Aeronautics and Astronautics, Aerospace Center, 370 L’Enfant Promenade, SW, Washington, DC 20024-2518, 1991. ISBN ?? LCCN ??

Berger:1998:AMR

[BL98] Marsha J. Berger and Randall J. LeVeque. Adaptive mesh refinement using wave-propagation algorithms for hyperbolic systems. SIAM Journal on Numerical Analysis, 35(6):2298–2316, December 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/
                                        article/31597.

Berger:2023:TAS

[BL23] Marsha J. Berger and Randall J. LeVeque. Towards adaptive simulations of dispersive tsunami propagation from an asteroid impact. In ICM—International Congress of Mathematicians. Vol.

7. Sections 15–20, pages 5056– 5071. EMS Press, Berlin, Germany, 2023. ISBN 3-9854706-5-0; 3-9854756-5-2; 3-9854705-8-8.

Berger:2024:IAM

[BL24] Marsha J. Berger and Randall J. LeVeque. Implicit adaptive mesh refinement for dispersive tsunami propagation. SIAM Journal on Scientific Computing, 46(4): B554–B578, ???? 2024. CO-DEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

                                        URL https://epubs.siam.org/
                                        doi/10.1137/23M1585210.

Berger:1994:FVM

[BLS94] M. J. Berger, R. J. LeVeque, and L. G. Stern. Finite volume methods for irregular onedimensional grids. In Gautschi [Gau94], pages 255–259. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Berger:1995:ATC

[BM95] Marsha Berger and John Melton.

An accuracy test of a Cartesian grid method for steady flow in complex geometries. RIACS Report 95-02, ????, ????, ???? 1995. Published in [BM96].

Berger:1996:ATC

[BM96] Marsha Berger and John Melton.

An accuracy test of a Cartesian grid method for steady flow in complex geometries. In J. Glimm et al., editors, Hyperbolic problems: theory, numerics, applications (Stony Brook, NY, 1994), pages 91–100. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1996. ISBN 981-02-2441-9. See technical report [BM95].

Berger:1993:ACG

[BME93] Marsha Berger, J. Melton, and F. Enomoto. 3D automatic Cartesian grid generation for Euler flows. In AIAA Paper 93-3386, 11th AIAA Computational Fluid Dynamics Conference, Orlando, Fla., July, 1993, page ?? American Institute of Aeronautics and Astronautics, Aerospace Center, 370 L’Enfant Promenade, SW, Washington, DC 20024-2518, 1993. ISBN ?? LCCN ??

Berger:1984:AMR

[BO84] Marsha Berger and Joseph Oliger. Adaptive mesh refinement for hyperbolic partial differential equations. Journal of computational physics, 53(3):484–512, March 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

Berger:1981:PCM

[BOR81] Marsha Berger, J. Oliger, and G. Rodrigue. Predictor-corrector methods for the solution of timedependent parabolic problems on parallel processors. In Schultz [Sch81], pages 197–201. ISBN 0-12-632620-7. LCCN QA377 .E53 1980. Lawrence Livermore Report UCRL-87522.

Berger:1992:APC

[BR92] Marsha Berger and Isidore Rigoutsos. An algorithm for point clustering and grid generation. IEEE Trans. Sys. Man. & Cyber., 21(5):1278–1286, September/October 1992. NYU Technical Report 501, April, 1990.

Berger:1993:SAM

[BS93] Marsha Berger and J. Saltzman. Structured adaptive mesh refinement on the Connection Machine. In Sincovec et al. [S^{+}93], page ?? ISBN 0-89871-315-3. LCCN QA76.58 .S55 1993 v.1-2. Two volumes.

Berger:1994:AC

[BS94] Marsha J. Berger and Jeff S. Saltzman. AMR on the CM-2. Applied Numerical Mathematics:

Transactions of IMACS, 14(1–3): 239–253, April 14, 1994. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

URL  http://www.elsevier.
com/cgi-bin/cas/tree/store/
apnum/cas_sub/browse/browse.
cgi?year=1994&volume=14&issue=

1-3&aid=452. Proceedings of the Third ARO Workshop on Adaptive Methods for Partial Differential Equations (Troy, NY, 1992). Also, RIACS Report No. 92-16.

Berger:1993:UMT

[BVM93] Marsha Berger, V. Venkatakrishnan, and D. Mavriplis. Unstructured multigrid through agglomeration. In Melson et al. [MMM93], page ?? ISBN ???? ISSN 0191-7811. LCCN ???? Two volumes.

Berger:2005:PSI

[BW05] Marsha Berger and Z. J. Wang.

Preface: [special issue in celebration of Antony Jameson’s 70th birthday]. International Journal of Computational Fluid Dynamics, 19(8):547, 2005. CODEN IJCFEC. ISSN 1061-8562 (print), 1029-0257 (electronic).

Delanaye:1999:AHC

[DAB^{+}99] M. Delanaye, M. Aftosmis, M. J.

Berger, Y. Liu, and T. Pulliam. Automatic hybrid-cartesian grid generation for high-Reynolds number flows around complex geometries. Technical Report AIAA Paper 99-0777, Reno, NV, American Institute of Aeronautics and Astronautics, 1801 Alexander Bell Drive, Suite 500, Reston, VA 20191, USA, January 1999.

DeGroot:1985:PIC

[DeG85] Doug DeGroot, editor. Proceedings of the 1985 International Conference on Parallel Processing, August 20–23, 1985. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1985. ISBN 0-8186-0637-1 (paperback), 0-8186-8637-5 (hardback), 0-8186-4637-3 (microfiche). LCCN QA76.6 .I613 1985; QA76.6 .I548 1985.

Dolling:1995:ACF

[DG95] David S. Dolling and Bernard Grossman, editors. 12th AIAA Computational fluid dynamics conference: a collection of technical papers: June 19–22, 1995, San Diego, CA. American Institute of Aeronautics and Astronautics, 1801 Alexander Bell Drive, Suite 500, Reston, VA 20191, USA, 1995. ISBN ???? LCCN ???? Two volumes.

Engquist:1990:TIC

[EG90] Bjorn Engquist and Bertil Gustafsson, editors. Third International Conference on Hyperbolic Problems: Theory, Numerical Methods and Applications: Proceedings, Uppsala, Sweden, June 11– 15, 1990. Studentlitteratur and Chartwell Bratt), Lund, Sweden and Bromley, Kent, England, 1990. ISBN 91-44-33421-4 (Studentlitteratur), 0-86238-285-8 (Chartwell Bratt). LCCN QA377.I563 1990. Two volumes.

Engquist:1985:LSC

[EOS85] Björn Engquist, Stanley Osher, and Richard C. J. Somerville, editors. Large-scale computations in fluid mechanics, volume 22 of Lectures in applied mathematics. American Mathematical Society, Providence, RI, USA, 1985. ISBN 0-8218-1122-3. xv + 370 (vol. 1), x + 409 (vol. 2) pp. LCCN QA901 .L37 1985. Two volumes.

Forrer:1999:FSC

[FB99] Hans Forrer and Marsha Berger.

Flow simulations on Cartesian grids involving complex moving geometries. In Hyperbolic problems: theory, numerics, applications, Vol. I (Zürich, 1998), volume 129 of Internat. Ser. Numer. Math., pages 315–324. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1999. ISBN 3-7643-6080-1.

Giuliani:2022:WSR

[GAB^{+}22] A. Giuliani, A. S. Almgren, J. B.

Bell, M. J. Berger, M. T. Henry de Frahan, and D. Rangarajan. A weighted state redistribution algorithm for embedded boundary grids. Journal of computational physics, 464(??):??, September 1, 2022. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0021999122003679.

Gautschi:1994:MCH

[Gau94] Walter Gautschi, editor. Mathematics of computation, 1943– 1993: a half-century of computational mathematics: Mathematics of Computation 50th Anniversary Symposium, August 9–13, 1993, Vancouver, British Columbia, volume 48 of Proceedings of Symposia in Applied Mathematics. American Mathematical Society, Providence, RI, USA, 1994. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Helzel:2014:CGE

[HB14] Christiane Helzel and Marsha J. Berger. Cartesian grid embedded boundary methods for hyperbolic problems. In Hyperbolic problems: theory, numerics, applications, volume 8 of AIMS Ser. Appl. Math., pages 675–683. Am. Inst. Math. Sci. (AIMS), Springfield, MO, USA, 2014. ISBN 1-60133-017-0, 1-60133-017-0.

Helzel:2005:HRR

[HBL05] Christiane Helzel, Marsha J. Berger, and Randall J. LeVeque. A high-resolution rotated grid method for conservation laws with embedded geometries. SIAM Journal on Scientific Computing, 26(3):785–809, May 2005. CO-DEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/43028.

Herbin:2002:FVC

[HK02] Raphaèle Herbin and Dietmar Kröner, editors. Finite volumes for complex applications III. International Symposium on Finite Volumes for Complex Applications (Third: 2002: Porquerolles Island, France). Kogan Page Science, London, UK, 2002. ISBN 1-903996-34-1. LCCN QA911 .I56 2002.

Hedstrom:1983:AMR

[HRBO83] Gerald Hedstrom, Garry Rodrigue, Marsha Berger, and Joseph Oliger. Adaptive mesh refinement for 1-dimensional gas dynamics. In Stepleman et al. [S^{+}83], pages 43–47. ISBN 0-444-86607-8. LCCN Q172 .I46 1982.

Krivodonova:2006:HOA

[KB06] Lilia Krivodonova and Marsha Berger. High-order accurate implementation of solid wall boundary conditions in curved geometries. Journal of computational physics, 211(2):492– 512, January 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0021999105002846.

Kutler:1997:FIC

[KFC97] Paul Kutler, J. (Jolen) Flores, and J. J. Chattot, editors. Fifteenth International Conference on Numerical Methods in Fluid Dynamics: proceedings of the conference held in Monterey, CA, USA, 24–28 June 1996, volume 490 of Lecture notes in physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 3-540-63054-6 (Berlin). ISSN 0075-8450 (print), 1616-6361 (electronic). LCCN QC150 .I59 1996.

LeVeque:2011:TMA

[LGB11] Randall J. LeVeque, David L. George, and Marsha J. Berger. Tsunami modelling with adaptively refined finite volume methods. Acta Numerica, 20:211–289, 2011. CODEN ANUMFU. ISBN ???? ISSN 0962-4929 (print), 1474-0508 (electronic).

Murman:2002:NSR

[MAB02] S. Murman, M. Aftosmis, and M. Berger. Numerical simulation of rolling-airframes using a multilevel Cartesian method. Technical Report AIAA Paper 2002-2798 (St. Louis, MO), American Institute of Aeronautics and Astronautics, 1801 Alexander Bell Drive, Suite 500, Reston, VA 20191, USA, June 2002. Published in [MAB04a].

Murman:2003:IAM

[MAB03a] Scott Murman, Michael J. Aftosmis, and Marsha J. Berger. Implicit approaches for moving boundaries in a 3-D Cartesian method. Technical Report AIAA Paper 2003-1119 (Reno, NV), American Institute of Aeronautics and Astronautics, 1801 Alexander Bell Drive, Suite 500, Reston, VA 20191, USA, January 2003.

Murman:2003:SDM

[MAB03b] Scott Murman, Michael J.

Aftosmis, and Marsha J. Berger. Simulations of 6-DOF motion with a Cartesian method. Technical Report AIAA Paper 2003-1246, American Institute of Aeronautics and Astronautics, 1801 Alexander Bell Drive, Suite 500, Reston, VA 20191, USA, January 2003. See also [MAB04b].

Murman:2004:NSR

[MAB04a] S. Murman, M. Aftosmis, and M. Berger. Numerical simulation of rolling-airframes using a multilevel Cartesian method. AIAA Journal of Spacecraft and Rockets, 41(3):??, ???? 2004. See also technical report [MAB02].

Murman:2004:SDM

[MAB04b] Scott Murman, Michael J.

Aftosmis, and Marsha J. Berger. Simulations of store separation from an F/A-18 with a Cartesian method. AIAA Journal of Aircraft, 41(4):??, ???? 2004. See also [MAB03b].

Mavriplis:2005:HRAa

[MAB05] Dimitri J. Mavriplis, Michael J.

Aftosmis, and Marsha Berger. High resolution aerospace applications using the NASA Columbia Supercomputer. In ACM [ACM05], page 61. ISBN 1-59593-061-2. LCCN ????

Mavriplis:2007:HRAb

[MAB07] Dimitri J. Mavriplis, Michael J.

Aftosmis, and Marsha Berger.

High resolution aerospace applications using the NASA Columbia Supercomputer. The International Journal of High Performance Computing Applications, 21(1):106–126, Spring 2007. CO-DEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic).

URL http://hpc.sagepub.com/
content/21/1/106.full.pdf+
html.

Mandli:2016:CBO

[MAB^{+}16] Kyle T. Mandli, Aron J. Ahmadia, Marsha Berger, Donna Calhoun, David L. George, Yiannis Hadjimichael, David I. Ketcheson, Grady I. Lemoine, and Randall J. LeVeque. Clawpack: building an open source ecosystem for solving hyperbolic PDEs. PeerJ Computer Science, 2:e68:1– e68:27, August 2016. ISSN 2376-5992.

May:2013:TDS

[MB13] Sandra May and Marsha Berger. Two-dimensional slope limiters for finite volume schemes on noncoordinate-aligned meshes. SIAM Journal on Scientific Computing, 35(5):A2163–A2187, ???? 2013. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

May:2014:MEI

[MB14] Sandra May and Marsha Berger. A mixed explicit implicit time stepping scheme for Cartesian embedded boundary meshes. In Finite volumes for complex applications VII. Methods and theoretical aspects, volume 77 of Springer Proc. Math. Stat., pages 393– 400. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. ISBN 3-319-05684-0, 3-319-05683-2.

May:2017:EIS

[MB17] Sandra May and Marsha Berger.

An explicit implicit scheme for cut cells in embedded boundary meshes. Journal of Scientific Computing, 71(3):919– 943, June 2017. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/
content/pdf/10.1007/s10915-
016-0326-2.pdf.

Melton:1995:ACG

[MBAW95] J. Melton, M. Berger, M. Aftosmis, and M. Wong. 3D applications of a Cartesian grid Euler method. Technical Report AIAA Paper 95-0853, American Institute of Aeronautics and Astronautics, 1801 Alexander Bell Drive, Suite 500, Reston, VA 20191, USA, January 1995.

Melson:1993:SCM

[MMM93] N. Duane Melson, T. A. Manteuffel, and S. F. McCormick, editors. Sixth Copper Mountain Conference on Multigrid Methods: proceedings of a workshop cosponsored by the National Aeronautics and Space Administration ... [et al.] and held at Copper Mountain, Colorado, April 4–9, 1993, number NSAS CP-3224 in NASA conference publication. National Aeronautics and Space Administration, Office of Management, Scientific and Technical Information Program, Washington, DC, USA, 1993. ISBN ???? ISSN 0191-7811. LCCN ???? Two volumes.

Roma:1999:AVI

[RPB99] Alexandre M. Roma, Charles S. Peskin, and Marsha J. Berger. An adaptive version of the immersed boundary method. Journal of computational physics, 153 (2):509–534, August 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

Stepleman:1983:SCA

[S^{+}83] R. S. (Robert S.) Stepleman et al., editors. Scientific computing: applications of mathematics and computing to the physical sciences, volume 1 of IMACS transactions on scientific computation. North-Holland, Amsterdam, The Netherlands, 1983. ISBN 0-444-86607-8. LCCN Q172 .I46 1982.

Sincovec:1993:PSS

[S^{+}93] Richard F. Sincovec et al., editors. Proceedings of the Sixth SIAM Conference on Parallel Processing for Scientific Computing, held March 22–24, 1993, in Norfolk, VA, USA. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1993. ISBN 0-89871-315-3. LCCN QA76.58 .S55 1993 v.1-2. Two volumes.

Soubbaramayer:1985:NIC

[SB85] Soubbaramayer and J. P. (Jean-Paul) Boujot, editors. Ninth International Conference on Numerical Methods in Fluid Dynamics: [proceedings], Gif-sur-Yvette/France 1984, volume 218 of Lecture notes in physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. ISBN 0-387-13917-6. LCCN QA911 .I54 1984; QA911 .I58 1984; TA357 .I564 1984. DM82.00 (West Germany).

Schultz:1981:EPS

[Sch81] Martin H. Schultz, editor. Elliptic problem solvers. Academic Press, New York, NY, USA, 1981. ISBN 0-12-632620-7. LCCN QA377 .E53 1980.

Sakell:1996:PFA

[SK96] L. Sakell and D. Knight, editors.

Proceedings of the First AFOSR Conference on Dynamic Motion CFD: New Brunswick, New Jersey, June 3–5, 1996. United States Air Force Office of Scientific Research, Washington, DC, USA, 1996. ISBN ???? LCCN ????

Wang:1996:TAG

[WB96] Z. J. Wang and M. J. Berger. Towards automatic grid independent viscous solutions with an adaptive Cartesian/quad grid flow solver. In Kutler et al. [KFC97], page ?? ISBN 3-540-63054-6 (Berlin). ISSN 0075-8450 (print), 1616-6361 (electronic). LCCN QC150 .I59 1996.

Wissink:2008:ACM

[WKB^{+}08] A. Wissink, B. Kosovic, M. Berger, K. Chand, and F. K. Chow. Adaptive Cartesian methods for modeling airborne dispersion. In Zomaya et al. [ZPCL08], page ?? ISBN ???? LCCN ????

Zomaya:2008:ACI

[ZPCL08] A. Zomaya, M. Parashar, S. Chandra, and X. Li, editors. Advanced Computational Infrastructures for Parallel and Distributed Adaptive Applications. Wiley Book Series on Parallel and Distributed Computing. Wiley, New York, NY, USA, 2008. ISBN ???? ???? pp. LCCN ????